OpenStudy (anonymous):

Mary is planning a circular rug. If she wants the rug to have a circumference of no more than 24 feet, which of the following represents the possible radius of her rug (to the nearest tenth foot)? Use 2 π r ≤ C. r ≤ 8.1 ft r ≤ 12 ft r ≤ 3.1 ft r ≤ 3.8 ft

OpenStudy (kropot72):

Rearrange the given equation to the following: \[r \le\frac{C}{2\pi}=\frac{24}{2\pi}\]